OpenStudy (pokemon23):

The larger of two numbers is 20 more than the smaller.Four times the larger is 70 more than 5 times the smaller.Find the numbers. The larger of two numbers is 20 more than the smaller.Four times the larger is 70 more than 5 times the smaller.Find the numbers. @Mathematics

OpenStudy (saifoo.khan):

Heyy

OpenStudy (saifoo.khan):

4(x+20) = 5*x + 70 4x+80 = 5x + 70 Subtract 4x from each side: 80=x+70 Subtract 70 from each side: 10=x x=10 Smaller x+20 = Larger 10+20= 30 Larger

OpenStudy (pokemon23):

ty my friend

OpenStudy (saifoo.khan):

uw